# Emergence of the Primacy Effect in Structured State-Space Models

### v2

@@ 2.2 Structured State-Space Models
For the polynomial coefficients to serve as a “memory” of the input signal, their temporal evolution must be trackable in an online manner; that is, the polynomial approximation at time $t$ should not refer back to past values of the input signal, $x|_{\leq t}(s)$ ($0\leq s<t$). Fortunately, for certain families of polynomials, including Legendre, Laguerre, and Fourier basis,222 The Fourier approximation (or transform) is not based on polynomials, but the theory can be generalized to incorporate it by taking the complex-valued basis $z^{n}:=e^{2\pi inx}$ and a measure on the unit circle (Gu et al.,, 2020). the coefficient dynamics can be described by an ordinary differential equation (ODE; Gu et al.,, 2020):
@@
where $\mathbf{h}(t):=(h_{1}(t),\dots,h_{N}(t))^{\mathsf{T}}$, and $A$ and $B$ are matrices of size $N\times N$ and $N\times 1$, respectively. The values on $A$ and $B$—referred to as the state and input matrices, respectively—depend on the choice of the underlying polynomial basis, and can also be adjusted via gradient-based optimization. A feedforward transformation of the state vector $\mathbf{h}(t)$ (achieved via left-multiplication by another matrix $C$) yields a (possibly multi-channel) output signal $\mathbf{y}(t)=C\mathbf{h}(t)\in\mathbb{R}^{M}$.333 The general formulation of the SSM incorporates an additional matrix $D$, which establishes a direct feedforward connection between the input and output signals, expressed as $\mathbf{y}(t)=C\mathbf{h}(t)+D\mathbf{x}(t)$. In practice, however, $D$ is often set to the identity matrix, effectively reducing the feedforward transformation to a simple residual connection (He et al.,, 2016). The resulting mapping $x\mapsto\mathbf{y}$ is termed the structured state-space model (hereinafter abbreviated as SSM; Gu et al.,, 2021; Gu et al., 2022b, ).
@@
Figure 1: Illustration of the continuous (left; $K(t):=e^{tA}B$) and discretized (right; $\bar{K}_{-j}:=\bar{A}^{j}\bar{B}$) SSM kernels based on Legendre polynomials with an exponentially decaying measure (Gu et al.,, 2023). The polynomial degree is indicated by the line colors. The kernels were evaluated at $t=-j\Delta t$ for $j=0,\dots,-255$ and $\Delta t\in\{0.001,0.01,0.1,1.0\}$. Increasing $\Delta t$ results in growing discrepancies between the continuous and discretized kernels.

### v3

@@ 2.2 Structured State-Space Models
For the polynomial coefficients to serve as a “memory” of the input signal, their temporal evolution must be trackable in an online manner; that is, the polynomial approximation at time $t$ should not refer back to past values of the input signal, $x|_{\leq t}(s)$ ($0\leq s<t$). Fortunately, for certain families of polynomials, including Legendre, Laguerre, and Fourier basis,222 The Fourier approximation (or transform) is not based on polynomials, but the theory can be generalized to incorporate it by taking the complex-valued basis $z^{n}:=e^{2\pi ins}$ and a measure on the unit circle (Gu et al.,, 2020). the coefficient dynamics can be described by an ordinary differential equation (ODE; Gu et al.,, 2020):
@@
where $\mathbf{h}(t):=(h_{0}(t),\dots,h_{N-1}(t))^{\mathsf{T}}$, and $A$ and $B$ are matrices of size $N\times N$ and $N\times 1$, respectively. The values on $A$ and $B$—referred to as the state and input matrices, respectively—depend on the choice of the underlying polynomial basis, and can also be adjusted via gradient-based optimization. A feedforward transformation of the state vector $\mathbf{h}(t)$ (achieved via left-multiplication by another matrix $C$) yields a (possibly multi-channel) output signal $\mathbf{y}(t)=C\mathbf{h}(t)\in\mathbb{R}^{M}$.333 The general formulation of the SSM incorporates an additional matrix $D$, which establishes a direct feedforward connection between the input and output signals, expressed as $\mathbf{y}(t)=C\mathbf{h}(t)+D\mathbf{x}(t)$. In practice, however, $D$ is often set to the identity matrix, effectively reducing the feedforward transformation to a simple residual connection (He et al.,, 2016). The resulting mapping $x\mapsto\mathbf{y}$ is termed the structured state-space model (hereinafter abbreviated as SSM; Gu et al.,, 2021; Gu et al., 2022b, ).
@@
Figure 1: Illustration of the continuous (left; $K(-t):=e^{tA}B$) and discretized (right; $\bar{K}_{-j}:=\bar{A}^{j}\bar{B}$) SSM kernels based on Legendre polynomials with an exponentially decaying measure (Gu et al.,, 2023). For a more intuitive visualization (particularly for readers unfamiliar with convolutional operations), the sign of the horizontal axis has been flipped (ranging in the negative domain) so that kernel values multiplied with past inputs appear on the left (unlike in the standard visualization of convolutional kernels, where they are placed on the right). The polynomial degree is indicated by the line colors. The kernels were evaluated at $t=j\Delta t$ for $j=0,\dots,255$ and $\Delta t\in\{0.001,0.01,0.1,1.0\}$. Increasing $\Delta t$ results in growing discrepancies between the continuous and discretized kernels.